# Genetic Data Governance in Crisis: Policy Recommendations for Safeguarding Privacy and Preventing Discrimination

### v1

@@
Abstract.
Genetic data collection has become ubiquitous today. The ability to meaningfully interpret genetic data has motivated its widespread use across forensics, clinical practice, and research, providing crucial insights into human health and ancestry while driving important public health initiatives. Easy access to genetic testing has fueled a rapid expansion of direct-to-consumer offerings, many of which are recreational in nature. However, the growth of genetic datasets and their applications has created significant privacy and discrimination risks, particularly as our understanding of the scientific basis for genetic traits continues to evolve. In this paper, we organize the uses of genetic data along four distinct ‘pillars’: clinical practice, research, forensic and government use, and recreational use. Using our scientific understanding of genetics, genetic inference methods and their associated risks, and existing regulatory mechanisms, we build a risk assessment framework that identifies key values that any governance system must preserve. We then analyze case studies from each of the pillars using this framework to assess how well existing legal and regulatory frameworks preserve desired values. Our investigation reveals critical gaps in existing regulatory frameworks and identifies specific threats to privacy and personal liberties, particularly through genetic discrimination. To address these challenges, we call for and propose comprehensive regulatory reforms including: (1) updating the legal definition of genetic data to protect against modern technological capabilities, (2) expanding the Genetic Information Nondiscrimination Act (GINA) to cover currently unprotected domains, and (3) establishing a unified regulatory framework under a single governing body to oversee all applications of genetic data.
We conclude with three open questions about genetic data: the challenges posed by its relational nature, including consent for relatives and minors; the complexities of international data transfer; and its potential integration into large language models.
Data Governance, Genetic Privacy, Regulatory Frameworks, Genetic Discrimination, Policy, Data Protection, Genetic Inference
$c$s: Applied computing Computational genomicsccs: Applied computing Health informaticsccs: Applied computing Lawccs: Applied computing Consumer health
1. Introduction
@@
The scale of genetic data generation has increased at a staggering rate since 2001 (Bick et al., 2024; Bycroft et al., 2018; Loos, 2020). A primary driver of this is decreased DNA sequencing costs (Wetterstrand, 2023), which have enabled the curation of large genetic datasets for personal and public health, ancestry inference, relative identification and advanced forensics. For example, there are whole genome sequences combined with medical records from approximately 400,000 US residents in the National Institutes of Health (NIH) All of Us Research Study, with plans to increase sampling to include at least 1 million US residents (National Institutes of Health, 2024). Direct-to-consumer (DTC) genetic testing companies have much more data: 23andMe has 14 million customer sequences (23a, 2023a) and Ancestry has 25 million (Ancestry, 2025). The increase in genetic data and improvements in technologies has resulted in the following aspirational “bold prediction” for 2030 by the NIH: “a person’s complete genome sequence along with informative annotations can be securely and readily accessible on their smartphone” (Green et al., 2020). From a scientific perspective, the collection of genetic data at this scale has been critically valuable to understanding health, ancestry, and driving public health initiatives.
@@
The propensity for genetic discrimination grows with poor protections for privacy. Since 2020, the U.S. Department of Homeland Security has accelerated its collection of genetic data, growing the CODIS database by more than 1.5 million people, the majority of whom are people of color (Glaberson et al., 2024). 23andMe—which boasts more than 14 million customers (23a, 2023a)—suffered a data breach after a hacker accessed 14,000 customer profiles [^1] While genetic data was not accessed, ancestry information and relative names were(23a, 2023b), primarily targeting individuals of Chinese and Ashkenazi Jewish descent (Carballo et al., 2024). These pressing, real-world discrimination and privacy concerns are exacerbated by the sordid history of eugenics and forced sterilization in U.S. states in the 20th century (a model later adopted by Nazi Germany (Spiegel, 2019)), and more recent practices of genetic determinism (Epstein, 2003), racial pseudoscience (Lala and Feldman, 2024; Duello et al., 2021; Harmon, 2018; Panofsky et al., 2021), race-motivated violence (Carlson et al., 2022), privacy and consent violations (Strand, 2016; Sterling, 2011; Jillson, 2024), and broad scientific distrust (Panacer, 2023; Kaye, 2012; Saulsberry and Terry, 2013). These are serious concerns that must be addressed today given the number of recent proposals for genetic data applications in public infrastructure, (e.g., to assess intelligence proxies, like IQ (R and S, 2018); insurance premiums (Karlsson Linnér and Koellinger, 2022); or educational outcomes (Harden et al., 2020)). The only federal-level legislation currently in effect is the Genetic Information Non-Discrimination Act (GINA) of 2008. GINA prohibits the use of genetic information in employment decisions and discrimination based on genetic information in health insurance coverage but does not prohibit genetic discrimination for long-term care, disability, or life insurance.
Our Work.
@@
In this paper, we make three contributions towards the construction of a robust genetic data governance system. First, we expand upon Wan et al.’s (Wan et al., 2022) Four Pillars of genetic data collectors and make the argument that inconsistent and leaky regulation across these pillars introduces opportunity for genetic discrimination and privacy violations. Second, we outline a risk assessment framework for modern-day genetic data governance and outline key values it should preserve. Third, we make three concrete recommendations to: (1) legally redefine genetic data; (2) expand GINA; and (3) create a regulatory framework for genetic data collectors.
Our paper is structured as follows: In Section 2, we provide an overview of genetic data and inference, with a focus on the nuances of genetic trait prediction. In Section 3, we introduce the extended Four Pillars and describe the existing legislative and regulatory structure governing genetic data use. In Section 4, we outline our risk assessment framework and apply it to four case studies in Section 5. We give our three policy recommendations in Section 6 and conclude, in Section 7, with three open challenges that are beyond the scope of this paper.
@@
2. Background: How meaning is derived from genetic data
@@
In this section, we provide a brief overview of genetic data and genetic inference. An individual’s genome is over 3 billion DNA bases [^2] Units of DNA long and derived equally from both genetic parents. Genetic variants—units of DNA that differ between individuals—are scattered throughout the genome and contain information about genetically influenced traits (from physical traits to behavioral traits). To predict traits or disease risk, geneticists analyze an individual’s unique set of genetic variants. However, this remains a challenging task: while diseases like cystic fibrosis or Huntington’s disease are determined by a single genetic variant, most traits are influenced by many variants across the genome as well as non-genetic factors (complex traits).
2.1. Your genetic data is shaped by your ancestors and shared with your relatives.
@@
The inheritance of DNA—where each parent passes down a random half of their genome to their offspring—results in relatives sharing portions of their DNA, with closer relatives sharing a larger proportion. This shared genetic material means that decisions made about any individual’s genome inherently extend to their genetic relatives. This relational nature of genetic data—the genetic dragnet—has been noted to significantly complicate discussions about privacy (Costello, 2022). A powerful example of the genetic dragnet came in 2018, when law enforcement leveraged it to identify the Golden State Killer. Authorities used DNA left at a crime scene to identify distant relatives through DNA matching on a public website, GEDMatch. By combining these matches with publicly available genealogical records, they identified the suspect despite the closest identifiable relatives sharing only a great-great-great-great-grandfather (Kaiser, 2018; Zabel, 2019). At the same time, an individual’s genome is unique to them: even the genomes of identical twins contain differences (Ormond et al., 2024). While a single genome in isolation may not always be immediately identifiable the relational nature of genetic data can enable re-identification. Researchers have argued that genomic data should be treated as ”always identifiable in principle” due to this inherent interconnectedness (Shabani and Marelli, 2019; Bonomi et al., 2020).
2.2. Geneticists use statistical models for trait prediction.
Inference using genetic data requires some knowledge of a variant’s effect on a trait or disease. This can sometimes be tested in animals by controlled experiments that manipulate the animal’s genome. However, this approach is infeasible and unethical to pursue in humans, and so researchers instead rely on large genetic datasets to ask ”Do individuals who have the variant tend to have the disease (or trait of interest) compared to those who do not have the variant?” To answer this question, researchers typically conduct genome-wide association studies (GWAS), which identify genetic variants that are associated with a trait and quantify the effect of each variant on the trait. The learned effect sizes of variants (also called weights) are directly interpretable [^3] For example, in a GWAS for height, an effect size of 0.01 would mean that the variant increases height by an average of 0.01 cm. and can be used for trait prediction. The polygenic score, considered the gold standard for trait prediction, is calculated as a weighted sum of genetic variants, with each variant scaled by its corresponding GWAS effect size. See Section 7 for a discussion of frontier models for genetic data (Fritzsche et al., 2023).
2.3. Insight from genetic data is a function of the cohort.
@@
The reference genetic dataset used for a GWAS is analogous to the training dataset in machine learning. Reference datasets contain genetic data and their “ground truth” annotations of trait values, whether it be ancestry, disease status, height, etc. Non-genetic factors of a cohort also influence a GWAS. Confounders like environmental exposures and social determinants of health can introduce bias in a genetic study. For example, the average participant in the UK Biobank  [^4] The UK Biobank is a leading resource which contains genetic and medical data from 500,000 UK residents aged 40-69 at the time of recruitment. tends to be more educated and healthier than the average Briton (Fry et al., 2017). The choice of the cohort—that is who is in the dataset and how their data is labeled555‘True” ancestry has been arbitrarily (and implicitly) set by the field to correspond to: where your ancestors lived 600 years ago (Coop, 2023). Individuals in the same ancestry group are distantly related in the same timescale (i.e., likely many ancestors in common dating back approximately 600 years ago). —is thus critical to a GWAS outcome. Prioritizing ”cohort matching” in a genetic study ensures that the test sample closely aligns with the reference cohort in key factors such as ancestry, age, geography, and social or environmental influences. This alignment minimizes confounding and enables a more accurate interpretation of genetic effects on traits. [^6] For example, (Mostafavi et al., 2020) find that genetic data prediction accuracy varies within an ancestrally-homogeneous white British cohort, e.g., weights from a BMI GWAS conducted in a younger cohort have higher prediction accuracy via a PGS in younger individuals than older individuals.
3. Background: Current Protections
@@
Figure 1. Four Pillars of genetic data collection and their regulatory considerations. Data collectors within each pillar operate with different collection strategies and intent to store, use, and transfer the data. However, there is only a patchwork of federal and state-level legislation that governs each Pillar. The resulting regulatory gaps lead to leaky protections through which genetic data—or knowledge derived from genetic data—can be used in other pillars with little to no oversight. Real-world examples [^8] Orchid Health is a private company that uses publicly available GWAS weights to compute polygenic risk scores (PGS) of embryos. Nucleus Genomics is also a company that similarly uses whole genome sequencing to compute PGS to predict health and behavioral traits. of leakage are indicated by red arrows.
@@
In this section, we outline the foundation of today’s genetic data collectors and the current regulatory and legislative environments that they are subject to. To better understand the vast landscape over which genetic data is being collected, we build upon the Four Pillars (Wan et al., 2022) of genetic data collectors. We categorize these data collectors by their motivations for collecting genetic data and their relationships with the individuals from whom the data is collected. We define the updated Four Pillars as the following:
@@
The “genetic data ecosystem” encompasses not only the entities within the Four Pillars but also the broader network of legislatures, regulators, and individuals involved in or affected by genetic data collection and use. By outlining this ecosystem, we aim to shed light on the regulatory gaps that enable genetic data leakage across pillars. We provide real-world examples of this leakage in Figure 8.
3.1. Federal protections
3.1.1. Genetic Information Non-Discrimination Act (GINA)
@@
GINA was enacted in 2008 and prohibits genetic discrimination by employers or health insurers. Specifically, GINA bars employers from making hiring, firing, or promotion decisions based on genetic information (and information disclosed about relatives) and bans health insurance providers from denying coverage or charging higher premiums based on genetic information including genetic predisposition for disease (Congress, 2008). GINA does not apply to life, disability, or long-term care insurance (Green et al., 2020) or any other venues outside of employment and health insurance (e.g., education). The US Equal Employment Opportunity Commission (EEOC) regulates workplace discrimination for employees and thus, is responsible for enforcing GINA in the workplace. Enforcement for health insurers falls under the purview of several federal agencies, including the Department of Health and Human Services (HHS); Department of Labor; Centers for Medicare and Medicaid Services; and the Department of the Treasury (Protections, OHRP).
3.1.2. Health Information Portability and Accountability Act (HIPAA)
@@
HIPAA, signed into law in 1996, grants security and privacy protections for patient health information (PHI). HIPAA applies to covered entities that maintain this information, such as healthcare and health insurance companies, but does not apply to many other entities that might retain health information, including search engines, medical information sites, or dating sites (Citron, 2020). Importantly, HIPAA does not apply to two of the largest DTC genetic testing companies 23andMe and AncestryDNA (Sklar, 2020). HIPAA does not cover deidentified data: PHI data with unique identifying information (names, address, phone numbers, dates, etc.) removed do not fall under HIPAA. Research consortia, biobanks, and research collaborations with DTC and clinical entities primarily use genetic data that is considered deidentified, and so HIPAA generally does not apply. However, the relational aspect of genetic data has motivated calls to “consider genomic data as, in principle, always identifiable” (Bonomi et al., 2020). The inability to completely strip data of all identifiable features introduces critical vulnerabilities for genome owners and exacerbates the risk of data leakage across the Four Pillars.
3.1.3. Food and Drug Administration (FDA)
@@
The FDA “is responsible for protecting the public health by ensuring the safety, efficacy, and security of human and veterinary drugs, biological products, and medical devices” (Commissioner, 2023). In the context of DTC genetic testing companies, the FDA only regulates well-defined high risk medical tests with clinical actions, such as 23andMe’s health tests for diseases like Parkinson’s, breast cancer, and late-onset Alzheimer’s. If the FDA approves a specific test, Centers for Medicare and Medicaid Services then choose to approve the test for insurance coverage (Daval and Kesselheim, 2023). Most genetic tests are not covered by Medicare, particularly DTC tests.
3.1.4. Federal Trade Commission (FTC)
@@
The FTC enforces federal laws that protect consumers from “fraud, deception and unfair business practices” (Commission, 2025). The FTC has warned DTC genetic testing companies that the results they return to customers must be backed by “reliable science”, and also warns companies against making exaggerated claims about the use of AI in their products (Jillson, 2024). Additionally, the FTC remains vigilant of deceptive practices regarding privacy policy changes and “dark patterns” designed to coerce consumers into consenting to data sharing [^9] In 2023, the Commission fined 1Health.io/Vitagene \$75,000, claiming they “left sensitive genetic and health data unsecured, deceived consumers about their ability to get their data deletes, and changed its privacy policy retroactively without adequately notifying and obtaining consent” (Commission, 2023).. The power and broad scope of the FTC—and their stated intentions of cracking down on DTC genetic testing companies (Jillson, 2024)—make it a major player in the regulatory space, and it will likely grow in importance if the federal government prioritizes genetic data privacy.
3.1.5. The Common Rule
@@
This is a federal policy that protects human subjects in research settings (Sciences et al., 2014). It was first codified in 1981 by the Department of Health, Education, and Welfare (now HHS) but adopted more widely by other federal agencies in 1991. The Common Rule institutionalizes protections for vulnerable participants, protections overseen by Institutional Review Boards (IRBs); IRBs are groups “formally designated to review and monitor biomedical research involving human subjects” (Research, 2024). Revisions enacted in 2017 particularly focus on informed consent for participants, but do not cover deidentified data (Menikoff et al., 2017; Protections, OHRP). IRBs ensure that consistent protection of human subjects are maintained and are necessary for research to begin and proceed if involving human participants. The maintenance of the Common Rule and HIPAA, in the case of medical data, is performed by IRBs.
3.2. State protections
@@
Figure 2. Beyond HIPAA and GINA, state-level genetic discrimination and privacy legislation varies widely in the United States. Left: States with genetic anti-discrimination legislation, distinguishing between states with both health insurance and employment protections (red) and those with health insurance protections only (orange). Right: states with general privacy laws that include genetic data (blue) and genetic privacy-specific laws (green).
@@
3.2.1. Genetic Anti-Discrimination Laws
@@
Most states have passed some version of a genetic anti-discrimination law (initially prompted by the passage of GINA) although the sector, scope, and strength of the laws differs between states. Some states offer fewer protections than GINA: Georgia, for example, has a statute outlawing genetic discrimination by health insurers, but not employers. Other states offer more protections than GINA. California’s CalGINA (Legislature, 2011) outlaws genetic discrimination in several sectors: employment, health insurance, housing, and by any state agency or entity receiving state funding (such as emergency medical services). Florida includes life, long-term care, and disability insurers (of Representatives, 2020). Many states (including Massachusetts and New Mexico) that include life insurance in their non-discrimination laws exempt companies if genetic data reliably, and based on “sound actuarial principles” (Legislature, 2025), gives information relating to mortality or morbidity (see Section 5.4 for a detailed case study).
3.2.2. Genetic Data Privacy Laws

### v2

# Principles and Policy Recommendations for Comprehensive Genetic Data Governance

@@
Abstract
Genetic data collection has become ubiquitous, producing genetic information about health, ancestry, and social traits. However, unregulated use—especially amid evolving scientific understanding—poses serious privacy and discrimination risks. These risks are intensified by advancing AI, particularly multi-modal systems integrating genetic, clinical, behavioral, and environmental data. In this work, we organize the uses of genetic data along four distinct ‘pillars’, and develop a risk assessment framework that identifies key values any governance system must preserve. In doing so, we draw on current privacy scholarship concerning contextual integrity, data relationality, and the Belmont principle. We apply the framework to four real-world case studies and identify critical gaps in existing regulatory frameworks and specific threats to privacy and personal liberties, particularly through genetic discrimination. Finally, we offer three policy recommendations for genetic data governance that safeguard individual rights in today’s under-regulated ecosystem of large-scale genetic data collection and usage.
Introduction
@@
The scale of genetic data generation has increased at a staggering rate since 2001 (Loos 2020; Bycroft et al. 2018). A primary driver of this is decreased DNA sequencing costs (Wetterstrand 2023). For example, there are whole genome sequences combined with medical records from approximately 400,000 US residents in the National Institutes of Health (NIH) All of Us Research Study, with plans to increase sampling to include at least 1 million US residents (National Institutes of Health 2024). Direct-to-consumer (DTC) genetic testing companies have much more data: 23andMe has 14 million customer sequences (23andMe 2023a) and Ancestry has 25 million (Ancestry 2025). The increase in genetic data and improvements in technologies has resulted in the following aspirational “bold prediction” for 2030 by the NIH: “a person’s complete genome sequence along with informative annotations can be securely and readily accessible on their smartphone” (Gunter and Green 2023). From a scientific perspective, the collection of genetic data and subsequent genetic information at this scale has been critically valuable to understanding health, ancestry, and driving public health initiatives.
@@
The unregulated collection of genetic data has created new avenues for surveillance, exploitation, and systemic harm. Since 2020, the U.S. Department of Homeland Security has accelerated its collection of genetic data, growing the federal DNA database (CODIS) by more than 1.5 million people, the majority of whom are people of color (Glaberson, Tse, and Tucker 2024). 23andMe suffered a data breach after a hacker accessed customer profiles, primarily targeting individuals of Chinese and Ashkenazi Jewish descent (Carballo, Schmall, and Tumin 2024). Such pressing, real-world discrimination and privacy concerns are exacerbated by the sordid history of eugenics and forced sterilization in U.S. states in the 20th century (a model later adopted by Nazi Germany (Spiegel 2019)), and more recent practices of genetic determinism (Epstein 2003), racial pseudoscience (Duello et al. 2021; Lala and Feldman 2024; Carlson et al. 2022; Panofsky, Dasgupta, and Iturriaga 2021), race-motivated violence (Carlson et al. 2022), privacy and consent violations (Jillson 2024; Sterling 2011; Strand 2016), and broad scientific distrust (Saulsberry and Terry 2013; Parkman et al. 2015; Kaye 2012). These serious concerns must be addressed today given the number of recent proposals to use genetic data and information in public infrastructure—e.g., to assess intelligence proxies, like IQ (Plomin and von Stumm 2018); insurance premiums (Karlsson Linnér and Koellinger 2022); or educational outcomes (Harden et al. 2020). The only federal-level legislation in effect is the Genetic Information Non-Discrimination Act (GINA), which prohibits using genetic information in employment and health insurance decisions but not in long-term care, disability, or life insurance.
@@
Our Work. In this paper, we make three contributions to construct a robust genetic data governance system. First, we expand upon Wan et al.’s (2022) Four Pillars of genetic data collectors and argue that inconsistent and leaky regulation across these pillars introduces opportunity for genetic discrimination and privacy violations. Second, we present a risk assessment framework for modern genetic data governance, derived from theoretical analysis and existing U.S. policy. Third, we make policy recommendations to (1) define genetic data and genetic information as a distinct protected category of sensitive data, (2) extend the Genetic Information Nondiscrimination Act (GINA) to include this category, and (3) establish a unified regulatory framework to govern all collection of genetic data and applications of genetic information.
@@
Background: How health and social information is derived from genetic data
@@
An individual’s genome is over 3 billion DNA bases long and derived equally from both genetic parents. Genetic variants—units of DNA that differ between individuals—are scattered throughout the genome and contain information about genetically influenced traits (from physical traits to behavioral traits). To predict traits or disease risk, geneticists analyze an individual’s unique set of genetic variants. However, this remains a challenging task: while diseases like cystic fibrosis or Huntington’s disease are determined by a single genetic variant, most traits are influenced by many variants across the genome as well as non-genetic factors (complex traits). In this section, we make the connection between genetic data—the raw genomic material—and genetic information, which refers to health, ancestry, and social traits inferred through statistical analysis of that genetic data.
Genetic data is inherently relational
The inheritance of DNA—where each parent passes down a random half of their genome to their offspring—results in relatives sharing portions of their DNA, with closer relatives sharing a larger proportion. Shared genetic material means that decisions made about any individual’s genome extend to their genetic relatives. At the population level, all individuals can be found in a latent web of interconnectedness. This means that a privacy breach in a small part of this web (that represents those in a genetic dataset, for example) may propagate to larger parts of the web and implicate individuals who were not in the original dataset. Thus, relationality applies to genetic data, where value is derived from shared connections among individuals (Viljoen 2024; Hargreaves 2017). This relationality complicates discussions of privacy in the context of genetic data (Costello 2022).
@@
A powerful example of the relationality of genetic data was highlighted in 2018, when law enforcement leveraged it to identify the Golden State Killer. Authorities used DNA left at a crime scene to identify distant relatives through DNA matching on a public website, GEDMatch. By combining these matches with publicly available genealogical records, they identified the suspect despite the closest identifiable relatives sharing only a great-great-great-great-grandfather (Kaiser 2018; Zabel 2019). An individual’s genome is shared with tens of thousands of other individuals, but at the same time is uniquely theirs: even the genomes of identical twins contain differences (Jonsson et al. 2021). While a single genome in isolation may not always be immediately identifiable, the relational nature of genetic data can enable re-identification. Researchers have argued that genomic data should be treated as “always identifiable in principle” due to this inherent interconnectedness (Bonomi, Huang, and Ohno-Machado 2020; Shabani and Marelli 2019).
Insight from genetic data is a function of the cohort
@@
Trait prediction models require a reference genetic dataset, which is analogous to the training dataset in machine learning. Reference datasets contain genetic data and their “ground truth” annotations of trait values, whether it be ancestry, disease status, height, etc. Confounders like environmental exposures and social determinants of health can introduce bias in a genetic study. For example, the average participant in the UK Biobank tends to be more educated and healthier than the average Briton (Fry et al. 2017). The choice of the cohort—that is who is in the dataset and how their data is labeled—is thus a critical consideration of trait prediction models and their generalizability, i.e., their out-of-distribution (OOD) performance in individuals outside of the cohort. Prioritizing “cohort matching” in a genetic study ensures that the test sample closely aligns with the reference cohort in key factors such as ancestry, age, geography, and social or environmental influences. This alignment minimizes confounding and enables a more accurate interpretation of strictly genetic effects on traits.
Trait prediction models from genetic data
Inference from genetic data requires some knowledge of a genetic variant’s effect on a trait or disease. This can sometimes be tested in animals by controlled experiments that manipulate the animal’s genome. However, this approach is infeasible and unethical to pursue in humans, and so researchers instead rely on large genetic datasets to ask “Do individuals who have the variant tend to have the disease (or trait of interest) compared to those who do not have the variant?” To answer this question, researchers typically conduct genome-wide association studies (GWAS), which identify genetic variants that are associated with a trait and quantify the effect of each variant on the trait. The learned effect sizes of variants (also called weights) are directly interpretable and can be used for trait prediction. The polygenic score (PGS), considered the gold standard for trait prediction, is calculated as a weighted sum of genetic variants, with each variant scaled by its corresponding GWAS effect size. GWAS/PGS approaches have been widely adopted for trait prediction.
The deployment of AI technologies for genetic inference is nascent but growing: predictive models are increasingly deployed to enable precision medicine therapies (Feiszt and Kenny 2025). Recently, there have been several calls to integrate genetic data with other information—rich clinical, behavioral, and environmental data using foundation models. Critically, biological data—particularly genomic sequences—is large-scale and shaped by noise, sparsity, and biased sampling across the diversity of life. In the domains of vision and language, integrating multi-modal data has shown success in improving predictive performance on sparse and noisy data domains. In the medical domain, Google’s Med-Gemini processes electronic health records (EHRs), imaging data (e.g., pathological, dermatological, and X-ray images), and genomic data (in the form of PGS) for trait prediction and other clinical uses (Yang et al. 2024). They claim Med-Gemini-polygenic—their model for predicting traits with PGS—to be highly performant on OOD data, a highly valued result in AI models for biology. However, it is important to consider that genetic data is often correlated with and confounded by environmental or non-biological variables. Thus, while OOD traits may appear to be genetically correlated with in-distribution (ID) traits (implying shared biological pathways), these correlations—particularly for behavioral phenotypes—are often confounded by non-genetic factors that mimic biological signal, undermining their interpretability as true genetic correlations (Border et al. 2022). Examples of this phenomenon have been demonstrated in places like genetic risk testing for opioid use disorder, where environmental covariates confounded genetic data (Pozzi 2023; Hatoum et al. 2021; Davis et al. 2025).
Blindspots for AI Applications Using Genetic Information
We believe that the increasing inclusion of genetic data into foundational models—while the interpretability of even simple linear models remains opaque—is of grave concern. GWAS are fraught with confounding: the effect sizes may be statistical artifacts due to correlations between genetic variants and environmental exposures; this is particularly true with behavioral traits. If variant-environmental correlations are not present in another population, the model will generalize poorly. Further, the non-zero effect size of the correlated variant will give the illusion of a genetic or biological basis for the trait, supporting ideas of genetic determinism. Statistical geneticists have developed study designs that control for almost all environmental confounding (Veller and Coop 2024; Howe et al. 2022; Nivard et al. 2022; Okbay et al. 2022), but these studies have not been widely adopted.
At the same time, researchers are using these existing studies—which suffer from confounding—to train AI models. We believe that the lack of consideration for this nuance in both training data curation and biological foundation model evaluations can exacerbate pre-existing issues in genetic data modeling approaches. While AI can improve trait prediction or interpretation, it may do so by modeling nonlinear interactions among correlated variables and leveraging gene–environment correlations. Yet, there remains an outsized excitement to develop increasingly complex AI models with genetic data (Yang et al. 2024). This can lead to compounded uncertainties due to the lack of interpretability and inherent complexity of AI models. We are particularly concerned about the use of multi-modal AI models incorporating genetic data to predict behavior, educational attainment, and other social outcome traits. PGS have already been developed for a variety of these traits, including educational attainment (Okbay et al. 2022), intelligence (Plomin and von Stumm 2018; Selzam S et al. 2019), standardized testing and mathematics scores (Harden et al. 2020; Smith-Woolley et al. 2018), substance use disorder (Barr et al. 2020), and aggression (Kretschmer et al. 2023). Med-Gemini-Polygenic was applied to health traits (i.e., not the aforementioned traits), and leveraged genetically correlated traits to perform well OOD (in this case, OOD refers to the trait, not the individuals). Genetically correlated health traits may reflect shared biology, but for behavioral and social traits, correlations can be induced by non-biological factors (Border et al. 2022) and thus an AI model may not be “learning” any biological insight. If even simple genetic models struggle with confounding from factors like environmental exposures and cohort biases, complex multi-modal AI models should not be assumed to inherently distinguish true genetic effects from these confounded associations, and may even leverage them, potentially exacerbating existing issues and compounding uncertainties.
Background: Current Protections
@@
Figure 1: The genetic ecosystem consists of the Four Pillars of genetic data collection and regulatory considerations. Data collectors within each pillar operate with different collection strategies and intent to store, use, and transfer the data. However, there is only a patchwork of federal and state-level legislation that governs each Pillar. The resulting regulatory gaps lead to leaky protections through which genetic data or genetic information can be used in other pillars with little to no oversight. Real-world examples of leakage are indicated by red arrows. Leakage refers to both (1) genetic data itself and (2) genetic information, i.e., inferences derived from genetic data.
@@
In this section, we outline the foundation of today’s genetic data collectors and the regulatory and legislative environments that they are subject to. To better understand the vast landscape over which genetic data is being collected, we build upon the Four Pillars (Wan et al. 2022) of genetic data collectors. We categorize these data collectors based on their motivations for collecting genetic data and the level of consent, control, and ownership afforded to the individuals from whom the data is collected. We define the updated Four Pillars as the following:
@@
The genetic data ecosystem encompasses not only the entities within the Four Pillars but also the broader network of legislatures, regulators, and individuals involved in or affected by genetic data collection and use. Here, we highlight the regulatory gaps that allow genetic data and genetic information to leak across pillars. We provide real-world examples of this leakage in Figure 1, illustrating the sociality problem formalized by Viljoen to contextualize how relational data’s social effects are categorized (Viljoen 2024).
Federal protections
Statute: Genetic Information Non-Discrimination Act (GINA).
@@
GINA was enacted in 2008 and prohibits genetic discrimination by employers or health insurers. Specifically, GINA bars employers from making hiring, firing, or promotion decisions based on genetic information (and information disclosed about relatives) and bans health insurance providers from denying coverage or charging higher premiums based on genetic information, including genetic predisposition for disease (Green, Lautenbach, and McGuire 2015). GINA does not apply to life, disability, or long-term care insurance (Gunter and Green 2023) or any other venues outside of employment and health insurance (e.g., education). The US Equal Employment Opportunity Commission (EEOC) regulates workplace discrimination for employees and thus, is responsible for enforcing GINA in the workplace. Before the passing of GINA, multiple cases of civil rights being violated in employment practices using genetic data have been recorded (Ajunwa 2016). Enforcement for health insurers falls under the purview of several federal agencies, including the Department of Health and Human Services (HHS); Department of Labor; Centers for Medicare and Medicaid Services; and the Department of the Treasury (Vermont Legislature 2019).
Statute: Health Information Portability and Accountability Act (HIPAA).
@@
HIPAA, signed into law in 1996, grants security and privacy protections for patient health information (PHI). HIPAA applies to covered entities that maintain this information, such as healthcare and health insurance companies, but does not apply to many other entities that might retain health information, including search engines, medical information sites, or dating sites (Citron 2020). Importantly, HIPAA does not apply to two of the largest DTC genetic testing companies 23andMe and AncestryDNA (Sklar 2020). HIPAA does not cover deidentified data: PHI data with unique identifying information (names, address, phone numbers, dates, etc.) removed do not fall under HIPAA. Research consortia, biobanks, and research collaborations with DTC and clinical entities primarily use genetic data that is considered deidentified, and so HIPAA generally does not apply. However, the relational aspect of genetic data has motivated calls to “consider genomic data as, in principle, always identifiable” (Bonomi, Huang, and Ohno-Machado 2020). The inability to completely strip data of all identifiable features introduces critical vulnerabilities for genome owners and exacerbates the risk of data leakage across the Four Pillars.
Agency: Food and Drug Administration (FDA).
@@
The FDA “is responsible for protecting the public health by ensuring the safety, efficacy, and security of human and veterinary drugs, biological products, and medical devices” (Commissioner 2023). In the context of DTC genetic testing companies, the FDA only regulates well-defined high risk medical tests with clinical actions, such as 23andMe’s health tests for diseases like Parkinson’s, breast cancer, and late-onset Alzheimer’s. If the FDA approves a specific test, Centers for Medicare and Medicaid Services then choose to approve the test for insurance coverage (Daval and Kesselheim 2023). Most genetic tests are not covered by Medicare, particularly DTC tests.
Agency: Federal Trade Commission (FTC).
@@
The FTC enforces federal laws that protect consumers from “fraud, deception and unfair business practices” (Commission 2024). The FTC has warned DTC genetic testing companies that the results they return to customers must be backed by “reliable science”, and also warns companies against making exaggerated claims about the use of AI in their products (Jillson 2024). Additionally, the FTC remains vigilant of deceptive practices regarding privacy policy changes and “dark patterns” designed to coerce consumers into consenting to data sharing. The power and broad scope of the FTC—and their stated intentions of cracking down on DTC genetic testing companies (Jillson 2024)—make it a major player in the regulatory space, and it will likely grow in importance if the federal government prioritizes genetic data privacy.
Policy: The Common Rule.
@@
This is a federal policy that protects human subjects in research settings (Sciences et al. 2014). It was first codified in 1981 by the Department of Health, Education, and Welfare (now HHS) but adopted more widely by other federal agencies in 1991. The Common Rule institutionalizes protections for vulnerable participants, protections overseen by Institutional Review Boards (IRBs); IRBs are groups “formally designated to review and monitor biomedical research involving human subjects” (Research 2019). Revisions enacted in 2017 focus on informed consent for participants, but do not cover deidentified data (Menikoff, Kaneshiro, and Pritchard 2017; Protections OHRP). IRBs ensure that consistent protection of human subjects are maintained and necessary for research to begin and proceed if involving human participants. The maintenance of the Common Rule and HIPAA, in the case of medical data, is performed by IRBs.
State protections
@@
Figure 2: State-level genetic discrimination and privacy legislation varies widely beyond federal HIPAA and GINA protections. Left: States with genetic anti-discrimination laws covering both health insurance and employment (red) versus health insurance only (orange). Right: States with general privacy laws including genetic data (blue) and genetic-specific privacy laws (green). States with both are green.
@@
Genetic Anti-Discrimination Laws.
@@
Most states have passed some version of a genetic anti-discrimination law (initially prompted by the passage of GINA) although the sector, scope, and strength of the laws differs between states. Some states offer fewer protections than GINA: Georgia, for example, has a statute outlawing genetic discrimination by health insurers, but not employers. Other states offer more protections than GINA. California’s CalGINA (Legislature 2011) outlaws genetic discrimination in several sectors: employment, health insurance, housing, and by any state agency or entity receiving state funding (such as emergency medical services). Florida includes life, long-term care, and disability insurers (of Representatives 2020). Many states (including Massachusetts and New Mexico) that include life insurance in their non-discrimination laws exempt companies if genetic data reliably, and based on “sound actuarial principles” (Legislature 2025), gives information relating to mortality or morbidity (see Section 5.4 for a detailed case study).
Genetic Data Privacy Laws.